OpenStudy (jared1):

Write the quadratic equation in general quadratic form below. 3x2 + 6x = 12

OpenStudy (t__prodigy):

Do u mean 3x^2+6x-12=0?????

OpenStudy (jared1):

yes

OpenStudy (t__prodigy):

Do u need the values of x? or what is your question exactly

OpenStudy (jared1):

it needs to be put in general quadratic form, and i have no clue on how to do that.

OpenStudy (jared1):

ax^2+bx+c=0 is the equation i think

OpenStudy (t__prodigy):

The general quadratic equation is ax^2 + bx+ c..>>invariably, the quadratic form of ur question is 3x^2+6x+(-12)=0. where a =3 b=6 and c=-12

OpenStudy (jared1):

thank you

OpenStudy (t__prodigy):

You welcome